# VoxATtack: A Multimodal Attack on Voice Anonymization Systems

### v1

@@ 1 Introduction
The aim of the VPAC is to develop attacker systems capable of undermining the privacy protections of anonymization frameworks. Specifically, participants are tasked with designing an automatic speaker verification (ASV) model which, given two anonymized speech utterances, determines whether they originate from the same speaker or not. Participants were instructed to develop their attacker systems against one or more anonymization systems chosen from three VPC 2024 baselines (B3, B4, B5) and four other top-performing participant-submitted systems (T8-5, T10-2, T12-5, T25-1) [2].
@@
Recently, the VPAC results have been announced, revealing the top-five performing attackers [7, 8, 9, 10, 11]. Among these, A.20 was the top-ranking attacker against all systems except T8-5 [7]. The authors achieved this via a multistage training strategy to finetune ResNet34 on ASV using low-rank adaptation, while enhancing its output by incorporating WavLM [12] features through the residual connections. Against T8-5, attacker A.5 [8] achieved the best results using ECAPA-TDNN and probabilistic linear discriminant analysis (PLDA) scoring, with contrastive loss regularization and data augmentation via SpecAugment [13] and the original (non-anonymized) speech. Similarly, [9] evaluated ECAPA-TDNN with multiple scoring metrics, e.g., PLDA and maximum similarity. However, they additionally applied voice normalization via kNN-VC [14] when attacking system T10-2. In [10], ECAPA-TDNN was also adopted and trained in two stages, with the feature extractor replaced by wav2vec $2.0$ outputs [15] and spectrogram resizing applied for data augmentation. In [11], TitaNet-Large [16] and ECAPA-TDNN were investigated under various training data conditions. Crucially, their results demonstrated that finetuning TitaNet-Large on all anonymized data simultaneously proved ineffective, suggesting the importance of considering the specific anonymization system when training an attacker. Finally, the work in [17], although not part of the challenge, showed the effectiveness of using phoneme durations for speaker de-anonymization.
@@ 2 Proposed Attacker Model
Audio Embedding Model ECAPA-TDNN [6] (Audio $\rightarrow$ $192$) (finetune) $22.2$ M
@@ 2.1 Model architecture
$\mathbf{w}_{\text{ens}}=[w_{f},w_{a},w_{t}]\in\mathbb{R}^{3}$ are the ensemble weights, with $w_{i}\in[0,1]$ and $\sum w_{i}=1$, and $\tilde{\mathbf{z}}\in\mathbb{R}^{C}$ denotes each classifier’s output, with $C$ denoting the number of training speakers. The ensemble weights are produced by a gating mechanism that takes as input the concatenation of $\mathbf{e}_{\text{audio}}$ and $\mathbf{e}_{\text{text}}$ along with their confidence scores $\gamma_{\text{audio}}$ and $\gamma_{\text{text}}$. This way, the model dynamically adapts the contribution of each modality to improve accuracy while avoiding overreliance on a single source of information.
@@ 2.2 Training
We finetune the ECAPA-TDNN parameters to adapt to anonymization distortions while keeping the BERT encoder weights fixed to preserve its pretrained linguistic knowledge. The rest of the model is trained from scratch. Our model is trained by optimizing a multi-task objective composed of four additive angular margin (AAM) loss terms [22], one for each of the audio, text, and fusion classifiers, plus one for the ensemble output. The training objective is formulated as:
@@ 3.2 Preprocessing
Speech segments were cropped to a maximum duration of $10$ seconds during training. The resulting segments were mean-centered and normalized to $-20$ dB root mean square level with maximum amplitude clipping at $1.0$. The corresponding transcriptions were left untruncated to preserve full contextual information. Although this introduces a mismatch in the duration of audio and text inputs, our architecture handles this gracefully by processing and aggregating each modality independently prior to fusion.
@@ 3.3 Implementation details
In our loss function 2, we use equal weighting for the auxiliary terms $\lambda_{f}=\lambda_{a}=\lambda_{t}=0.1$, and $\lambda_{e}=1$ for the primary ensemble loss. For the AAM parameters, we set the scale and margin to $30$ and $0.15$, respectively. We employed the AdamW optimizer [27] with a weight decay factor of $10^{-5}$. CyclicLR [28] was used to cycle the learning rate at every optimization step between $10^{-4}$ and $10^{-6}$, completing a full cycle every $13$k steps. Batch size was set to $32$ and the maximum number of training epochs was $25$, with an early stopping after patience of $10$ consecutive epochs without EER reduction on the validation set.
@@ 4.1 Attacker systems evaluation
The consistent improvements from $\text{ECAPA}_{\text{ours}}$ to VoxATtack demonstrates the effectiveness of our multimodal approach. The incorporation of textual information reduces EER by $0.7\%$ on B4 and up to $3.3\%$ on T8-5. These findings suggest exploitable intra-speaker linguistic patterns in the dataset, potentially exposing a vulnerability when using it in anonymization applications. To further validate our hypothesis, we fed only the transcriptions into the text branch and used $\mathbf{e}_{\text{text}}$ as speaker representation instead of $\mathbf{e}_{\text{fusion}}$. The resulting $\textrm{EER}_{\textrm{avg}}$ was $35.8\%$, which remarkably surpassed $\text{ECAPA}_{\text{baseline}}$ on all systems except B3, B4 and B5, despite the baseline access to the anonymized speech signals.
@@ 4.2 Cross-system evaluation
Finally, to inspect whether text fusion alone could explain the discrepancy with [11], we show the performance of $\text{ECAPA}_{\text{ours}}$ in the same settings. As shown in Figure 2, $\text{ECAPA}_{\text{ours}}$ achieved performance nearly identical to VoxATtack across all systems, with T8-5 being the only exception. These results show that the model has, in fact, improved relative to the single-system case in Section 4.1. This directly contradicts previous findings and warrants future investigation.
@@ Acknowledgment
Part of this work including sentences reformulation, code support, and Figures 2-3 were generated or assisted by Claude Sonnet 3.7 thinking and GPT-o1 models. Content was reviewed or edited by the authors.

### v2

@@ 1 Introduction
The aim of the VPAC is to develop attacker systems capable of undermining the privacy protections of anonymization frameworks. Specifically, participants are tasked with designing an automatic speaker verification (ASV) model which, given two anonymized speech utterances, determines whether they originate from the same speaker. Participants were instructed to develop their attacker systems against one or more anonymization systems chosen from three VPC 2024 baselines (B3, B4, B5) and four other top-performing participant-submitted systems (T8-5, T10-2, T12-5, T25-1) [2].
@@
Recently, the VPAC results have been announced, revealing the top-five performing attackers [7, 8, 9, 10, 11]. Among these, A.20 was the top-ranking attacker against all systems except T8-5 [7]. The authors achieved this via a multistage training strategy to fine-tune ResNet34 on ASV using low-rank adaptation, while enhancing its output by incorporating WavLM [12] features through the residual connections. Against T8-5, attacker A.5 [8] achieved the best results using ECAPA-TDNN and probabilistic linear discriminant analysis (PLDA) scoring, with contrastive loss regularization and data augmentation via SpecAugment [13] and the original (non-anonymized) speech. Similarly, [9] evaluated ECAPA-TDNN with multiple scoring metrics, e.g., PLDA and maximum similarity. However, they additionally applied voice normalization via kNN-VC [14] when attacking system T10-2. In [10], ECAPA-TDNN was also adopted and trained in two stages, with the feature extractor replaced by wav2vec $2.0$ outputs [15] and spectrogram resizing applied for data augmentation. In [11], TitaNet-Large [16] and ECAPA-TDNN were investigated under various training data conditions. Crucially, their results demonstrated that finetuning TitaNet-Large on all anonymized data simultaneously proved ineffective, suggesting the importance of considering the specific anonymization system when training an attacker. Finally, the work in [17], although not part of the challenge, showed the effectiveness of using phoneme durations for speaker de-anonymization.
@@ 2 Proposed Attacker Model
Audio Embedding Model ECAPA-TDNN [6] (Audio $\rightarrow$ $192$) (fine-tune) $22.2$ M
@@ 2.1 Model architecture
$\mathbf{w}_{\text{ens}}=[w_{f},w_{a},w_{t}]\in\mathbb{R}^{3}$ are the ensemble weights, with $w_{i}\in[0,1]$ and $\sum_{i}w_{i}=1$, and $\tilde{\mathbf{z}}\in\mathbb{R}^{C}$ denotes each classifier’s output, with $C$ denoting the number of training speakers. The ensemble weights are produced by a gating mechanism that takes as input the concatenation of $\mathbf{e}_{\text{audio}}$ and $\mathbf{e}_{\text{text}}$ along with their confidence scores $\gamma_{\text{audio}}$ and $\gamma_{\text{text}}$. This way, the model dynamically adapts the contribution of each modality to improve accuracy while avoiding overreliance on a single source of information.
@@ 2.2 Training
We fine-tune the ECAPA-TDNN parameters to adapt to anonymization distortions while keeping the BERT encoder weights fixed to preserve its pretrained linguistic knowledge. The rest of the model is trained from scratch. Our model is trained by optimizing a multi-task objective composed of four additive angular margin (AAM) loss terms [22], one for each of the audio, text, and fusion classifiers, plus one for the ensemble output. The training objective is formulated as:
@@ 3.2 Preprocessing
Speech segments were cropped to a maximum duration of $10$ seconds during training. The resulting segments were mean-centered and normalized to $-20$ dB root mean square level, with maximum amplitude clipping at $1.0$. The corresponding transcriptions were left untruncated to preserve full contextual information. Although this introduces a mismatch in the duration of audio and text inputs, our architecture handles this gracefully by processing and aggregating each modality independently prior to fusion.
@@ 3.3 Implementation details
In our loss function 2, we use equal weighting for the auxiliary terms $\lambda_{f}=\lambda_{a}=\lambda_{t}=0.1$, and $\lambda_{e}=1$ for the primary ensemble loss. For the AAM parameters, we set the scale and margin to $30$ and $0.15$, respectively. We employed the AdamW optimizer [27] with a weight decay factor of $10^{-5}$. CyclicLR [28] was used to cycle the learning rate at every optimization step between $10^{-4}$ and $10^{-6}$, completing a full cycle every $13$k steps. The batch size was set to $32$ and the maximum number of training epochs was $25$, with an early stopping after $10$ consecutive epochs without EER reduction on the validation set.
@@ 4.1 Attacker systems evaluation
The consistent improvements from $\text{ECAPA}_{\text{ours}}$ to VoxATtack demonstrate the effectiveness of our multimodal approach. The incorporation of textual information reduces EER by $0.7\%$ on B4 and up to $3.3\%$ on T8-5. These findings suggest exploitable intra-speaker linguistic patterns in the dataset, potentially exposing a vulnerability when using it in anonymization applications. To further validate our hypothesis, we fed only the transcriptions into the text branch and used $\mathbf{e}_{\text{text}}$ as speaker representation instead of $\mathbf{e}_{\text{fusion}}$. The resulting $\textrm{EER}_{\textrm{avg}}$ was $35.8\%$, which remarkably surpassed $\text{ECAPA}_{\text{baseline}}$ on all systems except B3, B4 and B5, despite the baseline having access to the anonymized speech signals.
@@ 4.2 Cross-system evaluation
Finally, to inspect whether text fusion alone could explain the discrepancy with [11], we analyzed the performance of $\text{ECAPA}_{\text{ours}}$ in the same settings. As shown in Figure 2, $\text{ECAPA}_{\text{ours}}$ achieved performance nearly identical to VoxATtack across all systems, with T8-5 being the only exception. These results show that the model has, in fact, improved relative to the single-system case in Section 4.1. This directly contradicts previous findings and warrants future investigation.
@@ Acknowledgment
Part of this work, including sentence reformulation, code support, and Figures 2-3, was generated or assisted by Claude Sonnet 3.7 thinking and GPT-o1 models. Content was reviewed or edited by the authors.